OpenStudy (anonymous):

The sum of two consecutive integers is 203. What are the integers?

OpenStudy (sburchette):

We can let one integer equal x. The next integer (since they are consecutive) will by x+1. Now we set up the equation.\[x+(x+1)=203\]\[2x+1=203\]\[2x=202\]\[x=101\]So if the first integer is 101, then the next integer will be 102. We can test this to see if our values are correct. 101+102=203.

OpenStudy (anonymous):

So what would the answer be?

OpenStudy (sburchette):

The two integers are 101 and 102.

OpenStudy (anonymous):

I just need one answer